40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 20 14:42:34 2011, steve, Update, SUS, oplev gains today 
    Reply  Wed Jul 20 19:24:12 2011, rana, Update, SUS, oplev gains today Untitled.pnga.png
       Reply  Wed Jul 20 19:48:03 2011, Jamie, Update, SUS, Re: oplev gains today 
       Reply  Thu Jul 21 09:04:59 2011, valera, steve, Update, SUS, oplev gains were not optimized 
       Reply  Mon Jul 25 00:35:44 2011, rana, Update, SUS, something wrong with ETMY LR sensor a.pngbad.png
          Reply  Tue Jul 26 11:38:45 2011, steve, Update, SUS, ETMY bias checked 
    Reply  Thu Jul 21 16:06:35 2011, steve, Update, SUS, oplev gains today checked by conlog 
Message ID: 5014     Entry time: Thu Jul 21 16:06:35 2011     In reply to: 5001
Author: steve 
Type: Update 
Category: SUS 
Subject: oplev gains today checked by conlog 

 Present values agree with conlog records. It can be concluded that there were no big changes made. There are some 0.1-0.2 gain  and one polarity changes during the periods of Valera's visits.

 

ELOG V3.1.3-